# HAFO: A Force-Adaptive Control Framework for Humanoid Robots in Intense Interaction Environments

### v3

@@ 3.3 Upper-body motion generation based on imitation learning
$q_{\text{curr }}^{u}(t)$ $=q_{0}(t)+\alpha_{i}\left(q_{\text{tar }}^{u}(t)-q_{0}(t)\right)$ (2)
Among them, $q_{0}(t)$ represents the actual joint angle vector at time t, $q_{\text{tar}}(t)$ is the target joint angle vector of the upper-body at that time $t$, and $\alpha_{i}\in[0,1]$ is the course gain coefficient, which modulates the motion amplitude of the upper-body reference trajectory.
We adopt an asymmetric actor-critic training approach, where the critic networks use both observable information $o_{t}$ and privileged information $o_{o}^{\text{priv}}=\left\{F_{\text{ee}},F_{\text{torso}},v_{\text{lin}}\right\}$ to achieve accurate value estimation, while the actor networks rely solely on the observable information of the physical machine.
@@ 3.4 Virtual spring-damper system
In complex and unpredictable environments, a precise and stable control system is crucial for achieving task objectives. To improve the robustness of humanoid robots in the presence of unknown external disturbances, this paper introduces a pulling-force modeling mechanism based on spring-damper system [44, 45, 46], explicitly incorporating the pulling force interaction process into the training loop, aiming to regulate the robot’s body dynamics through a virtual spring-damper model, the corresponding dynamic model is given by
$\mathrm{m}\ddot{\mathrm{x}}=K_{\mathrm{p}}\left(\mathrm{x}_{\text{des }}-\mathrm{x}\right)+K_{\mathrm{d}}\left(\dot{\mathrm{x}}_{\text{des }}-\dot{\mathrm{x}}\right)+\mathrm{f}_{\text{ext }}$ (3)
Where $K_{p}$ and $K_{d}$ represent the stiffness and damping of the virtual spring-damper system, $\left(\mathrm{x}_{\mathrm{des}},\dot{\mathrm{x}}_{\mathrm{des}}\right)$ denotes the desired position and velocity, and $f_{spring}$ represents the spring-damper force.
[…]
@@ 4 Simulation Experiments
In this section, we conduct a series of force-interaction simulation experiments and comprehensively evaluated the effectiveness of the proposed HAFO framework through quantitative comparisons with the baseline methods. All simulations are conducted in MuJoCo via zero-shot transfer from Isaac Gym. The cross-platform validation approach effectively mitigates overfitting risks inherent to a single simulator, thereby effectively verifying policy robustness.
@@ 4.1 Force adaptive control under hand-induced disturbances
[…]
@@
HAFO 0.22±0.04 0.41±0.03 0.46±0.04 0.33±0.02 0.48±0.03 0.53±0.05
Furthermore, ablation experiments are conducted to quantify the contribution of each HAFO module. We sequentially ablate (i) the dual-agent architecture, retaining only the lower-body agent (w.o. DA), and (ii) the force-adaptative module (w.o. Force). As shown in Table 1, comparative experiments demonstrate that HAFO achieves superior tracking performance, both in terms of velocity tracking and upper-body angle following. The Upper-FIX-Lower-RL policy simplifies training by fixing the upper-body and relying exclusively on lower-body control for locomotion. However, it fails to guarantee system stability when the upper-body executes dynamic motions in real-world scenarios. Notably, HAFO demonstrates a more pronounced performance advantage as the hand-applied external force increases. Furthermore, ablation results demonstrate that removing either module leads to performance degradation. The random upper-body motion introduced during training serves as an adversarial perturbation source, potentially compelling the lower-body policy to learn more generalized dynamic compensation mechanisms. Meanwhile, force-interaction curriculum learning effectively enhances the dual-agent’s robustness and adaptability under perturbations.
@@ 4.2 Force adaptive control under rope suspension
Table 2: The comparison of different methods in terms of action smoothness and upper-body motion tracking error.
@@
Upper-OL-Lower-RL 2.49±2.38 0.27±0.11
Upper-FIX-Lower-RL 5.66±4.50 0.29±0.15
HAFO w.o. DA 1.60±0.49 0.27±0.24
HAFO -w/o- force 4.52±1.77 1.16±0.66
HAFO(Ours) 0.38±0.16 0.20±0.04
@@ 4.3 Force-Adaptive Control under Thrust Disturbance
Figure 3: Policy performance under various external thrust disturbances. (a) Performance under continuous constant-force disturbances applied from multiple directions. (b) Performance under one-second transient force disturbances from multiple directions.
To systematically assess policy robustness against thrust perturbations, we conduct two disturbance rejection experiments: applying sustained constant forces and one-second transient forces along eight directions in the horizontal plane, progressively increasing force magnitude until loss of stability. As shown in Figure 3, HAFO demonstrates greater tolerance to external disturbances in most directions compared to HAFO w.o. DA, HAFO w.o. Force, and the baseline strategy, indicating that adversarial upper-body training and the force curriculum both improve the robot’s force adaptation under thrust perturbations.
@@ 5 Real-world experiments
We deployed the HAFO strategy on physical Unitree G1 humanoid robot and conduct a series of hardware experiments. As shown in Figure 4, the HAFO controller completes multiple force interaction tasks, collectively validating the technical performance and practical applicability of our method on the physical robot.
Figure 4: HAFO enables force-adaptive control for humanoid robots across multiple scenarios. (a) Whole-body control with sandbag loads. (b) Locomotion while carrying a box with one hand. (c) Being lifted onto an elevated platform and resuming locomotion. (d) Locomotion while carrying flower pots with both hands. (e) Policy deployment in a suspended state. (f) Performing dance movements. (g) Window cleaning while suspended by ropes.
@@
[…]
@@
HAFO 0.16±0.05 0.40±0.04 0.39±0.03 0.41±0.03
5.1 Loco-manipulation under hand loads
@@
[…]
@@
5.2 Stable control under suspension
Traditional policy initialization schemes typically require robots to sequentially execute the series of procedures: foot-ground contact, posture adjustment, policy execution, and rope release. This process is relatively cumbersome, and sustained external disturbances from the rope readily leads to robot instability. This experiment validates the feasibility of suspended startup. The HAFO controller can initiate directly from a rope-suspended state and rapidly reach steady state, significantly simplifying the policy deployment process. Humanoid robots are typically controlled based on high-level signals such as target velocity and foot contact commands. Under the dynamic interference of rope tension, significant changes occur in the robot’s locomotion states, including linear velocity, angular velocity, body posture, and foot-ground contact. The baseline strategy misinterprets the normal swinging of the rope as robot instability, triggering emergency actions that cause violent oscillations. In contrast, the HAFO policy leverages environmental feedback for adaptive compensation, robustly maintaining locomotion stability.
@@ 6 Conclusion
In this paper, we propose HAFO, a force-adaptive control framework for humanoid robots. HAFO employs a dual-agent control strategy and explicitly introduces various dynamic disturbances, achieving stable motion control and precise motion imitation. During the model training, a spring-damper system is established, and fine-grained tension control of the pulling force is achieved through manipulation of virtual springs. Additionally, HAFO employs asymmetric Actor-Critic architecture wherein the critic-network is granted full access to privileged state variables, thereby furnishing a more accurate approximation of state-action value. Experimental results demonstrate that the proposed strategy enables the humanoid robots to exhibit strong adaptability under challenging conditions such as heavy load, thrust disturbance and rope suspension, providing robust support for solving substantial force interaction tasks in real-world applications.
Appendix A RL training
We use Proximal Policy Optimization (PPO) [41] as the base algorithm, and Table A.1 presents the detailed training hyperparameters.
Table 1: RL Controller Hyperparameters.
Hyperparameter Value
Clip param 0.2
Discount factor($\gamma$) 0.99
Value loss coef 1.0
Entropy coef 0.01
Actor learning rate 0.001
Critic learning rate 0.001
Weight decay 0.01
Max grad norm 1.0
Desired KL 0.01
Num steps per env 24
MLP size [512, 256, 128]
Appendix B State and Action Space
This section provides a detailed description of the state and action spaces for the dual-agent policy. The state space for both the upper and lower body strategies is identical, consisting of the whole-body state set. The dimensions of the action space for the upper and lower body strategies correspond to the number of degrees of freedom (DOF) in the humanoid robot’s upper and lower bodies, respectively. Specific details are shown in Table B.2.
@@ Appendix B State and Action Space
Commands 5 (Vel+Stand) 14 (Motion angle) 19
@@
Actions 12 (Lower) 15 (Upper) 27 (Whole)
Total dim 81 93 110
@@
Appendix C Reward Design
This paper adopts a reward function design similar to that in [2, 6, 19], while making targeted additions and modifications to certain reward components to better accommodate the needs of full-body control tasks and force adaptation control. These adjustments aim to enhance both the training effectiveness and the robustness of the strategy. The specific reward components are detailed in Table C.3.
@@ Appendix D Domain randomization
To enhance the robustness of the policy and reduce the performance gap between the simulation and real environments, we applied domain randomization to key parameters in the simulation. This approach allows the policy better adapt to unknown factors when transferred to the real environment. The specific domain randomization parameters and their ranges are listed in Table D.4.
@@ Appendix E Extend HAFO to Large-size Humanoid Robots
As shown in Figure E.1, We further extend the HAFO strategy to the full-scale humanoid robot H1-2 (height: 1.78 m, weight: 70 kg). Due to its larger mass and limited actuators, this robot presents greater challenges in terms of control. However, with minimal parameter adjustments, stable control is achieved within the simulation environment, thereby validating the strategy’s strong generalization and scalability.
Figure 1: Multi-scenario validation of the HAFO policy on the unitree H1-2 humanoid robot. (a) Robot upper-body swing. (b) Punching and maintaining stable standing. (c) Stable locomotion with a 100 N external force applied to each hand. (d) Stable control in the suspended state. (e) Rapid robot adjustment after external impact.

### v4

@@ 3.3 Upper-body motion generation under external forces
$q_{\text{curr }}^{u}(t)$ $=q_{0}^{u}(t)+\alpha_{i}\left(q_{\text{tar }}^{u}(t)-q_{0}^{u}(t)\right)$ (2)
Among them, $q_{0}^{u}(t)$ represents the actual joint angle vector at time $t$, $q_{\text{tar}}^{u}(t)$ is the target joint angle vector of the upper-body at that time $t$, and $\alpha_{i}\in[0,1]$ is the course gain coefficient, which modulates the motion amplitude of the upper-body reference trajectory.
We adopt a constrained residual action space, in which the policy does not directly learn absolute joint targets $\theta_{t}^{\text{target}}$, but instead outputs a corrective offset $\mathrm{a}_{t}^{u}$ relative to the sampled upper-body reference trajectory joint positions $\theta_{t}^{\text{ref}}$. To prevent training collapse caused by adversarial interactions between the dual agents, the output of $\mathrm{a}_{t}^{u}$ is constrained, ensuring that the network’s action outputs remain within a controllable range.
$\theta_{t}^{\mathrm{target}}=\theta_{t}^{\mathrm{ref}}+\mathrm{a}_{t}^{u},\ \text{s.t. }|\mathrm{a}_{t}^{u}|\leq\Delta$ (3)
@@ 3.4 Virtual spring-damper system
In complex and unpredictable environments, a precise and stable control system is crucial for achieving task objectives. To improve the robustness of humanoid robots in the presence of unknown external disturbances, this paper introduces a tension-force modeling mechanism based on spring-damper system [14, 15, 33], explicitly incorporating the tension force interaction process into the training loop, aiming to regulate the robot’s body dynamics through a virtual spring-damper model, the corresponding dynamic model is given by
$m\ddot{x}=K_{\mathrm{p}}\left(x_{\text{des }}-x\right)+K_{\mathrm{d}}\left(\dot{x}_{\text{des }}-\dot{x}\right)+f_{\text{ext }}$ (4)
Where $K_{p}$ and $K_{d}$ represent the stiffness and damping of the virtual spring-damper system, $\left(x_{\mathrm{des}},\dot{x}_{\mathrm{des}}\right)$ denotes the desired position and velocity, and $f_{\text{ext }}$ represents the external force.
Figure 2: Spring-damper model and performance analysis. (a) Spring-damper model schematic on the humanoid robot. (b)Pelvis position, spring force, and ground-reaction force versus a specified $x_{\mathrm{des}}^{\text{pelvis }}$ displacement for Spring-damper and stiffness models.
Leveraging a spring-damper model, we accurately capture the mechanics of external tensile force. By progressively adjusting the $x_{\mathrm{des}}^{\text{pelvis }}$, the tensile force is tuned to achieve stable position control, as shown in Fig. 2. Comparative tests show that stiffness control produces pronounced oscillations relative to spring-damper model, confirming the critical role of the damping term in enhancing stability and precision. We integrate the spring-damper system into the training loop, applying the force application position, cycle, and dropout timing randomization to effectively enhance the policy’s generalization capability.
Figure 3: Unitree G1 Humanoid robot sim2sim results. We evaluate the model’s performance in diverse intense-interaction environments. (a) Achieving stable control in a suspension scenario. (b) Achieving a smooth, autonomous transition from a suspended state to a ground movement state. (c) Swinging both arms while each hand bears a 30 N external force. (d) Punching while each hand bears a 30 N external force. (e) Rapidly recovering balance after experiencing a lateral 250 N thrust disturbance lasting 0.2 seconds. (f) Rapidly recovering balance after experiencing a longitudinal 250 N thrust disturbance lasting 0.2 seconds.
@@ 4 Simulation Experiments
In this section, we conduct a series of force-interaction simulation experiments and comprehensively evaluated the effectiveness of the proposed HAFO framework through quantitative comparisons with the baseline methods. All simulations are conducted in MuJoCo [38] via zero-shot transfer from Isaac Gym [26]. The cross-platform validation approach effectively mitigates overfitting risks inherent to a single simulator, thereby effectively verifying policy robustness. As shown in Fig. 3, we display the evaluation results of HAFO across various force disturbance scenarios, providing comprehensive validation of the policy’s force adaptability.
@@ 4.1 Force adaptive control under hand-induced disturbances
Methods $\mathrm{E}_{\text{tracking}}^{\text{upper}}\downarrow$ $\mathrm{E}_{\text{tracking}}^{\text{root}}\downarrow$
@@
HAFO(Ours) 0.22±0.04 0.41±0.03 0.46±0.04 0.33±0.02 0.48±0.03 0.53±0.05
Furthermore, ablation experiments are conducted to quantify the contribution of each HAFO module. We sequentially ablate (i) the dual-agent architecture, retaining only the lower-body agent (w.o. DA), and (ii) the force-adaptative module (w.o. Force). As shown in Table 1, comparative experiments demonstrate that HAFO achieves superior tracking performance, both in terms of velocity tracking and upper-body angle following. The upper-FIX-lower-RL policy simplifies training by fixing the upper-body and relying exclusively on lower-body control for locomotion. However, it fails to guarantee system stability when the upper-body executes dynamic motions. Notably, HAFO demonstrates a more pronounced performance advantage as the hand-applied external force increases. Furthermore, ablation results demonstrate that removing either module leads to performance degradation. The random upper-body motion introduced during training serves as an adversarial perturbation source, potentially compelling the lower-body policy to learn more generalized dynamic compensation mechanisms. Meanwhile, force-interaction curriculum learning effectively enhances the dual-agent’s robustness and adaptability under perturbations.
@@ 4.2 Force adaptive control under rope suspension
Table 2: Comparison of different methods in terms of action smoothness and upper-body motion tracking error.
   Methods    $\Delta a\downarrow$    $E_{\text{tracking}}^{\text{upper}}\downarrow$
@@
[…]
Figure 4: Policy performance under external thrust disturbances. (a) Performance under continuous constant-force disturbances applied from multiple directions. (b) Performance under one-second transient force disturbances from multiple directions.
@@ 4.3 Force-Adaptive Control under Thrust Disturbance
To systematically assess policy robustness against thrust perturbations, we conduct two disturbance rejection experiments: applying sustained constant forces and one-second transient forces along eight directions in the horizontal plane, progressively increasing force magnitude until loss of stability. As shown in Fig. 4, HAFO demonstrates greater tolerance to external disturbances in most directions compared to HAFO w.o. force, and the baseline strategy, indicating that adversarial upper-body training and the force curriculum both improve the robot’s force adaptation under thrust perturbations.
@@ 5 Real-world experiments
We deployed HAFO strategy on physical Unitree G1 humanoid robot and conduct a series of hardware experiments. As shown in Fig. 5, the HAFO controller completes multiple force interaction tasks, collectively validating the technical performance and practical applicability of our method on the physical robot.
5.1 Loco-manipulation under hand loads
Figure 5: HAFO enables force-adaptive control for humanoid robots across multiple scenarios. (a) Achieving whole-body control with sandbag loads. (b) Walking while carrying a box with one hand. (c) Being lifted onto an elevated platform and resuming locomotion. (d) Walking while carrying flower pots with both hands. (e) Deploying the policy in a suspended state. (f) Performing dance movements. (g) Cleaning windows while being suspended by ropes.
@@ 5.1 Loco-manipulation under hand loads
$\mathrm{E}_{\text{tracking}}^{\text{upper}}\downarrow$ $\mathrm{E}_{\text{tracking}}^{\text{root}}\downarrow$ $\mathrm{E}_{\text{tracking}}^{\text{upper}}\downarrow$ $\mathrm{E}_{\text{tracking}}^{\text{root}}\downarrow$
@@
HAFO(Ours) 0.16±0.05 0.40±0.04 0.39±0.03 0.41±0.03
5.2 Stable control under suspension
Traditional policy initialization schemes typically require robots to sequentially execute the series of procedures: foot-ground contact, posture adjustment, policy execution, and rope release. This process is relatively cumbersome, and sustained external disturbances from the rope readily leads to robot instability. This experiment validates the feasibility of suspended startup. The HAFO controller can initiate directly from a rope-suspended state and rapidly reach steady state, significantly simplifying the policy deployment process. Humanoid robots are typically controlled based on high-level signals such as target velocity and foot contact commands. Under the dynamic interference of rope tension, significant changes occur in the robot’s locomotion states, including linear velocity, angular velocity, body posture, and foot-ground contact, as shown in Fig. 6. The baseline strategy misinterprets the normal swinging of the rope as robot instability, triggering emergency actions that cause violent oscillations. In contrast, the HAFO policy leverages environmental feedback for adaptive compensation, robustly maintaining locomotion stability.
@@ 5.2 Stable control under suspension
   Tasks    $\Delta a\downarrow$    $\mathrm{E}_{\text{tracking}}^{\text{upper}}\downarrow$
@@
Figure 6: Motion analysis of the robot under rope suspension conditions. (a) Under the influence of rope swing, the robot experiences dynamic changes in linear velocity $v_{\text{lin}}$, angular velocity $\omega_{\text{yaw}}$, and body tilt angle $\theta$, and other motion parameters during the suspension process. (b) The HAFO strategy integrates environmental feedback, enabling the velocity-commanded motion policy to achieve adaptive control in the suspended state.
@@ 6 Conclusion
In this paper, we propose HAFO, a force-adaptive control framework for humanoid robots. HAFO employs a dual-agent control strategy and explicitly introduces various dynamic disturbances, achieving stable motion control and precise motion imitation. During the model training, a spring-damper system is established, and fine-grained tension control of the tension force is achieved through manipulation of virtual springs. Additionally, HAFO employs asymmetric Actor-Critic architecture wherein the critic-network is granted full access to privileged state variables, thereby furnishing a more accurate approximation of state-action value. Experimental results demonstrate that the proposed strategy enables the humanoid robots to exhibit strong adaptability under challenging conditions such as heavy load, thrust disturbance and rope suspension using a single dual-agent policy, providing robust support for solving substantial force interaction tasks in real-world applications.
Appendix A State and Action Space
This section provides a detailed description of the state and action spaces for the dual-agent policy. The state space for both the upper and lower body strategies is identical, consisting of the whole-body state set. The dimensions of the action space for the upper and lower body strategies correspond to the number of degrees of freedom (DOF) in the humanoid robot’s upper and lower bodies, respectively. Specific details are shown in Table 1.
@@ Appendix A State and Action Space
Commands 4 (Vel cmd + Stand cmd) 14 (Motion angles) 18
@@
Actions 15 (Lower) 14 (Upper) 29 (Whole)
Total dim 83 92 111
@@
Appendix B Reward Design
This paper adopts a reward function design similar to that in [13, 48, 5], while making targeted additions and modifications to certain reward components to better accommodate the needs of full-body control tasks and force adaptation control. These adjustments aim to enhance both the training effectiveness and the robustness of the strategy. The specific reward components are detailed in Table 2.
@@ Appendix B Reward Design
Appendix C RL training
We use Proximal Policy Optimization (PPO) [32] as the base algorithm, and Table 3 presents the detailed training hyperparameters.
Table 3: RL controller hyperparameters.
    Hyperparameter     Value
    Clip param     0.2
    Discount factor ($\gamma$)     0.99
    Value loss coef     1.0
    Entropy coef     0.01
    Actor learning rate     0.001
    Critic learning rate     0.001
    Weight decay     0.01
    Max grad norm     1.0
    Desired KL     0.01
    Num steps per env     24
    MLP size     [512, 256, 128]
@@ Appendix D Domain randomization
To enhance the robustness of the policy and reduce the performance gap between the simulation and real environments, we applied domain randomization to key parameters in the simulation. This approach allows the policy better adapt to unknown factors when transferred to the real environment. The specific domain randomization parameters and their ranges are listed in Table 4.
@@ Appendix E Extend HAFO to Large-size Humanoid Robots
As shown in Fig. 1, we further extend the HAFO strategy to the full-scale humanoid robot H1-2 (height: 1.78 m, weight: 70 kg). The increased mass and limited actuators make controlling the robot more challenging. However, with minimal parameter adjustments, stable control is achieved within the simulation environment, thereby validating the strategy’s strong generalization and scalability.
Figure 1: Multi-scenario validation of the HAFO policy on the unitree H1-2 humanoid robot. (a) Swinging the robot’s upper body. (b) Balancing under Punching motion. (c) Walking with a 100 N external force applied to each hand. (d) Achieving stable control in the suspended state. (e) Rapidly recovering balance after being subjected to a thrust disturbance.